# Robust Preference Optimization through Reward Model Distillation

### v1

@@ 3 Uncertainty-aware reward model distillation
As discussed in the previous section, a core issue in preference optimization is that the true preference distribution $p^{*}(y_{1}\succ y_{2}\mid x)$ is not known. Attempting to infer it from finite-sample preference data (that may further be biased or out-of-distribution with respect to the target domain) can then result in a failure to learn reasonable policies. In this section, we now propose an inherently regularized approach to direct preference optimization that uses uncertainty-aware reward model distillation.
@@ 3.1 Reward model distillation
Suppose for the moment that the reward function $r^{*}$ was in fact known, and did not have to be inferred from sampled preference data. Under this setting, we can then define an efficient offline optimization procedure that is similar in spirit to DPO, but no longer relies directly on a preference dataset. Concretely, given unlabeled samples $(x,y_{1},y_{2})\sim\rho$ (where the number of samples can be potentially unlimited), we can define a simple “distillation” loss, $\mathcal{L}_{\mathrm{distill}}(r^{*},\pi_{\theta})$, as follows:
@@
Intuitively, the distillation loss seeks to exactly match differences in reward model scores across all generation pairs $(x,y_{1},y_{2})$. It is then easy to see that under the Bradley-Terry model, this is equivalent to matching the strength of the preference relationship, $y_{1}\succ y_{2}$. Furthermore, by only matching differences, we can still conveniently ignore the log partition term, $\log Z(x)$, in the implicit reward formulation for $\pi_{\theta}$ as shown in (4), as it is constant across different $y$ for any given $x$. Finally, similar to the motivation in DPO, we can show that minimizing $\mathcal{L}_{\mathrm{distill}}(r^{*},\pi_{\theta};\rho)$ indeed results in an optimally aligned policy $\pi_{\theta^{*}}$, as long as the data distribution $\rho$ has sufficient support.
@@
The above result holds for a broad class of data distributions $\rho(x,y_{1},y_{2})$, and makes no assumptions on $r^{*}$ (e.g., it is no longer necessary for it to be defined using a Bradley-Terry model). In fact, this result can also be seen as strict generalization of the IPO framework of [3] when taking $r^{*}(x,y)\triangleq\mathbf{1}\{y=y_{w}\}$, if labeled pairs $(x,y_{w},y_{l})$ are provided instead of the unlabeled pairs $(x,y_{1},y_{2})$.
Of course, the true reward $r^{*}$ is usually not known in practice. Still, as in standard RLHF, we can go about constructing good proxies by using the preference data to identify plausible target reward models $r_{\mathrm{tgt}}$—further guided by any amount of regularization and inductive bias that we desire. A natural choice is to first learn $r_{\mathrm{tgt}}$ on the preference data $\mathcal{D}_{\mathrm{pref}}$ using standard methods, and then reuse $\mathcal{D}_{\mathrm{pref}}$ to distill $\pi_{\theta}$, which is similar to classical settings in teacher-based model distillation [14; 26]. Furthermore, as $r_{\mathrm{tgt}}$ is a real-valued model, at a bare minimum it is guaranteed to induce a regularized Bradley-Terry preference distribution $p_{\mathrm{tgt}}(y_{1}\succ y_{2}\mid x)>0,~{}\forall x,y_{1},y_{2}\in\mathcal{X}\times\mathcal{Y}\times\mathcal{Y}$, and thereby avoid some of the degeneracies identified in §2.3 for the maximum likelihood estimate under DPO.
@@ 3.2 Pessimistic reward model distillation
$\max_{\pi_{\theta}}\underset{r_{\mathrm{tgt}}^{i}\in\mathcal{S}}{\min}\mathbb{E}_{\mu(x)}\Big{[}\underbrace{\mathbb{E}_{\pi_{\theta}(y\mid x)}[r_{\mathrm{tgt}}^{i}(x,y)]-\mathbb{E}_{\pi_{\mathrm{ref}}(y\mid x)}[r_{\mathrm{tgt}}^{i}(x,y)]}_{\text{ advantage over baseline policy}}-\beta\mathbb{D}_{\mathrm{KL}}(\pi_{\theta}(\cdot\mid x)\|\pi_{\mathrm{ref}}(\cdot\mid x))\Big{]}.$ (8)
@@
In this pessimistic objective we are no longer optimizing $\pi_{\theta}$ for a single reward, but optimizing $\pi_{\theta}$ to produce generations that are scored favorably on average, even by the worst-case reward model in the set $\mathcal{S}$, relative to the generations of the baseline policy $\pi_{\mathrm{ref}}$. [^1] It is useful to optimize the advantage as it cancels the effects of constant differences between reward models in $\mathcal{S}$. We are also free to use any baseline policy $\pi_{\mathrm{base}}$; we pick $\pi_{\mathrm{ref}}$ for simplicity and ease of analysis in §5. When the set $\mathcal{S}=\{r^{*}\}$ consists of only the ground-truth reward, the objective (8) is equivalent to standard RLHF (1), up to a constant offset independent of $\theta$. More generally, whenever $\mathcal{S}$ includes a good proxy $\widetilde{r}$ for $r^{*}$, the pessimistic advantage evaluation ensures that the the policy $\pi^{*}_{\theta}$ that maximizes eq. 8 still has a large advantage over $\pi_{\mathrm{ref}}$ under all $r\in\mathcal{S}$, including $\widetilde{r}$. This use of pessimism to handle uncertainty in the knowledge of the true reward is related to similar techniques in the offline RL literature [16; 5].
For the objective to be meaningful, the set $\mathcal{S}$ has to be chosen carefully. When $\mathcal{S}$ is small, it might not include any good proxy for $r^{*}$. Conversely, if $\mathcal{S}$ is too rich, it forces $\pi_{\theta^{*}}$ to be nearly identical to $\pi_{\mathrm{ref}}$, since any deviations from $\pi_{\mathrm{ref}}$ might be penalized by some reward model in $\mathcal{S}$. Consequently, we want to design $\mathcal{S}$ to be the smallest possible set which contains a reasonable approximation to $r^{*}$.
To optimize (8), it turns out that we can formulate it as an equivalent constrained offline optimization problem, that we will show to conveniently admit a similar loss form as (7).
@@ 3.2.1 Optimization
The constraint in (9) can then be relaxed and approximately optimized by introducing an objective with a Lagrangian-style penalty with strength $\alpha>0$ on a form of distillation loss as (7), i.e.,
@@
Assume the same conditions as Theorem 1. Then for any $0<\gamma<\infty$, there exists a $\lambda\geq 0$ such that $\pi_{\theta^{*}}(y\mid x)\in\arg\!\min_{\pi_{\theta}}\mathcal{L}_{\mathrm{pdistill}}(\mathcal{S},\pi_{\theta};\rho)$, where $\pi_{\theta^{*}}$ is a minimizer over all possible policies, is a solution to (9) for the effective reward model set
@@ 3.3 Pessimistic DPO
We can also observe that Proposition 2 can be leveraged to obtain an alternative, implicitly pessimistic, objective that uses DPO directly instead of distillation. Consider the following regularized DPO loss:
@@ 4 Experimental results
The main motivation for reward distillation and pessimism is to increase alignment robustness in challenging settings where it is difficult to learn good policies directly from the preference data. To demonstrate the effectiveness of our approach, we run experiments on the popular TL;DR summarization task [29; 32], in which we simulate a scenario where the preference data has a spurious correlation between the length of a summary and whether or not it is preferred. [^4] Length has been repeatedly shown in the past to correlate with reward [28; 21].
@@ 4.1 Experimental setup
We first train an “oracle” reward model on the TL;DR preference data training set [29] and relabel all preference pairs with this oracle. This enables us to use the oracle reward model for evaluation, without worrying about the gap to true human preferences. After relabeling, longer responses (where longer is defined as $y_{1}$ having at least $10\%$ more tokens than $y_{2}$) are preferred in $61\%$ of the examples.
To test the effect of a spurious correlation on preference-based policy optimization, we select as a training set 30K examples from the relabeled data such that the longer output is preferred in $\rho$ fraction of examples, with $\rho\in\{0.2,0.3,0.4,0.5,0.6,0.7,0.8\}.$ Each such training set is denoted $\mathcal{D}_{\rho}$. At each $\mathcal{D}_{\rho}$, we compare our approach to DPO [23] and IPO [3], which are currently the most commonly used offline alignment methods. We test the following variants of distillation and pessimism:
@@
  • Pessimistic Distilled DPO (pd-DPO): Combines the above two by training a reward model $r_{\rho}$ on $\mathcal{D}_{\rho}$ and optimizing the pessimistic distillation objective (Eq. (11)) with confidence set $\mathcal{S}=\{r_{\mathrm{tgt}}\}$.
  • Pessimistic Ensemble DPO (e-DPO): To create ensembles of reward models, we subsample from each $\mathcal{D}_{\rho}$ five preference datasets, $\mathcal{D}_{\rho,b}$, at $b\in\mathcal{B}=\{0.2,0.4,0.5,0.6,0.8\}$, such that the fraction of pairs where the longer response is preferred is $b$, and train reward models $r_{\rho,b}$ on those subsets. Consequently, sensitivity to length should vary across ensemble members. We then apply the same procedure as pd-DPO above, with a confidence set $\mathcal{S}_{\rho}=\{r_{\rho,b}\}_{b=1}^{\mathcal{B}}$.
@@ 4.2 Results
Figure 2: Main results, showing the advantage in oracle reward compared to the initial finetuned policy. Errorbars correspond to bootstrap 95% confidence intervals for finite sample variance. Ensemble DPO (e-DPO) is significantly better than DPO and IPO in the challenging setup where shorter responses are preferred ($\rho\leq 0.5$), and is generally the best-performing method overall in this regime. Distilled DPO (d-DPO) performs best when longer responses are preferred ($\rho>0.6$).
We present the results of our experiment in Figure 2. As can be seen in the plot, the more challenging setting is when $\rho<0.5$, which corresponds to a sample of preference annotations in which shorter outputs are generally preferred. This distribution shift is more difficult because as mentioned the oracle reward model (trained on human annotations) has a bias in favor of longer outputs [28]. Nevertheless we get sizable improvements compared to the reference policy $\pi_{\textrm{SFT}}$ for all length bias values.
@@ 5 Theoretical analysis
This section characterizes problems with the DPO objective and solutions offered by pessimistic DPO and distillation, focusing on the simplified scenario in which we optimize with respect to a single preference pairs $(y^{w},y^{\ell})$. Once again, all proofs are deferred to Appendix A.
In its Lagrangian formulation, pessimistic DPO adds a forward KL term to the DPO objective (§3.3). For the sake of analysis, we assume that the preference annotations are sampled from the reference distribution, $\mu(x)\times\pi_{\mathrm{ref}}(y\mid x)\times\pi_{\mathrm{ref}}(y\mid x)$. Then a finite-sample approximation of the forward KL term is $\hat{\Omega}(\Theta):=\sum_{(y^{w},y^{\ell})\in\mathcal{D}_{\mathrm{Pref}}}-(\log\pi_{\theta}(y^{\ell})+\log\pi_{\theta}(y^{w})).$ By applying this finite-sample approximation, p-DPO has a finite optimum, unlike DPO, as shown in Proposition 1. Note that this analysis is limited in two ways: (1) as mentioned, we compute the KL term over the completions in the preference data; (2) we directly optimize the probability ratios $\psi_{w}=\pi_{\theta}(y^{w})/\pi_{\mathrm{ref}}(y^{w})$ and $\psi_{\ell}=\pi_{\theta}(y^{\ell})/\pi_{\mathrm{ref}}(y^{\ell})$, rather than optimizing them jointly through the parameters. For sufficiently expressive $\pi_{\theta}$, however, this approximation captures the behavior of the two algorithms reasonably well.

### v2

@@ 5 Uncertainty-aware reward model distillation
As discussed in the previous section, a core issue in preference optimization is that the true preference distribution $p^{*}(y_{1}\succ y_{2}\mid x)$ is not known. Attempting to infer it from finite-sample preference data (which may further be biased or out-of-distribution with respect to the target domain) can then result in a failure to learn reasonable policies. In this section, we propose a regularized, pessimistic approach to direct preference optimization that brings explicit reward modeling back into the picture through a model distillation objective, while still maintaining the simplicity and efficiency of offline alignment methods.
@@ 5.1 Reward model distillation
Suppose for the moment that the reward function $r^{*}$ was in fact known, and did not have to be inferred from sampled preference data. Under this setting, we can then construct a straightforward and efficient offline optimization procedure that is similar in spirit to DPO, but no longer relies directly on a preference dataset. Concretely, given unlabeled samples $(x,y_{1},y_{2})\sim\rho$ (where the number of samples can be potentially unlimited), we can define a simple squared “distillation” loss that matches the pairwise differences of the explicit reward $r^{*}$ with those of the implicit policy reward defined by $\pi_{\theta}$, i.e.,
@@
Due to symmetry, here it does not matter if $(y_{1},y_{2})=(y^{w},y^{\ell})$, i.e., preferred vs. dispreferred, or vice versa. Notably, a similar squared-loss objective has also been recently proposed and shown to be effective in independent work by Mao et al. (2024) and Gao et al. (2024b). We now first provide further motivation for this approach, before extending it to pessimistic variants in the next section.
Intuitively, the distillation loss in (7) seeks to exactly match differences in reward model scores across all generation pairs $(x,y_{1},y_{2})$. It is easy to see that under the Bradley-Terry model, this is equivalent to matching the strength of the preference relationship, $y_{1}\succ y_{2}$. Furthermore, by only matching differences, we can still conveniently ignore the log partition term, $\log Z(x)$, in the implicit reward formulation for $\pi_{\theta}$ as shown in (4), as it is constant across different $y$ for any given $x$. Finally, similar to the motivation in DPO, we can show that minimizing $\mathcal{L}_{\mathrm{distill}}(r^{*},\pi_{\theta};\rho)$ indeed results in an optimally aligned policy $\pi_{\theta^{*}}$, as long as the data distribution $\rho$ has sufficient support over the space of prompts and responses.
@@
The theorem holds for a broad class of data distributions $\rho(x,y_{1},y_{2})$, and makes no assumptions on $r^{*}$. For example, it is no longer necessary for it to be defined using a Bradley-Terry model. Notably, it applies for reward models that are much larger and potentially better than the policy model, as the reward model is not used at test time (in contrast to DPO, which ties the size of the policy model used for generation to the size of the implicit reward model that is trained on preferences). In fact, this result can also be seen as strict generalization of the IPO framework of Azar et al. (2024), which corresponds to the special case $r^{*}(x,y)\triangleq\mathbf{1}\{y=y_{w}\}$, if labeled pairs $(x,y_{w},y_{l})$ are provided instead of the unlabeled pairs $(x,y_{1},y_{2})$.
Of course, the true reward $r^{*}$ is usually not known in practice. Still, as in standard RLHF, we can construct good proxies by using the preference data to identify plausible target reward models $r_{\mathrm{tgt}}$, further guided by any amount of regularization and inductive bias that we desire. Moreover, prior work has found that such explicitly trained reward models are more accurate and generalize better than the implicit reward model defined in (4) that is learned by DPO Tang et al. (2024a); Lin et al. (2024). A natural choice is to thus first learn $r_{\mathrm{tgt}}$ on the preference data $\mathcal{D}_{\mathrm{pref}}$ using standard methods, and then reuse $\mathcal{D}_{\mathrm{pref}}$ to distill $\pi_{\theta}$, which is similar to classical settings in teacher-based model distillation Hinton et al. (2015); Romero et al. (2015). Furthermore, as $r_{\mathrm{tgt}}$ is a real-valued model, at a bare minimum it is guaranteed to induce a regularized Bradley-Terry preference distribution $p_{\mathrm{tgt}}(y_{1}\succ y_{2}\mid x)>0,~{}\forall x,y_{1},y_{2}\in\mathcal{X}\times\mathcal{Y}\times\mathcal{Y}$, and thereby avoid the degeneracies identified in §4 for the maximum likelihood estimate under DPO.
Takeaways: Reward model distillation If the true reward $r^{*}$ is known, Eq. (7) provides an objective that distills $r^{*}$ directly into $\pi_{\theta^{*}}$, where $\pi_{\theta^{*}}$ maximizes $r^{*}$ as in Eq. (1). When the true reward $r^{*}$ is unknown, we can use any form of approximate reward model $r_{\mathrm{tgt}}$. Furthermore, any real-valued $r_{\mathrm{tgt}}$ naturally adds regularization, and avoids degenerate optima such as the 00-probability solutions identified in §4.
@@ 5.2 Pessimistic reward model distillation
$\max_{\pi_{\theta}}\underset{r_{\mathrm{tgt}}^{i}\in\mathcal{S}}{\min}\mathbb{E}_{\mu(x)}\Big{[}\underbrace{\mathbb{E}_{\pi_{\theta}(y\mid x)}[r_{\mathrm{tgt}}^{i}(x,y)]-\mathbb{E}_{\pi_{\mathrm{ref}}(y\mid x)}[r_{\mathrm{tgt}}^{i}(x,y)]}_{\text{ advantage over the baseline policy}}-\beta\mathbb{D}_{\mathrm{KL}}(\pi_{\theta}(\cdot\mid x)\|\pi_{\mathrm{ref}}(\cdot\mid x))\Big{]}.$ (8)
@@
In this pessimistic objective we are no longer optimizing $\pi_{\theta}$ for a single reward, but optimizing $\pi_{\theta}$ to produce generations that are scored favorably on average, even by the worst-case reward model in the set $\mathcal{S}$, relative to the generations of the baseline policy $\pi_{\mathrm{ref}}$. [^3] It is useful to optimize the advantage as it cancels the effects of constant differences between reward models in $\mathcal{S}$. We are also free to use any baseline policy $\pi_{\mathrm{base}}$; we pick $\pi_{\mathrm{ref}}$ for simplicity and ease of analysis in §7. When the set $\mathcal{S}=\{r^{*}\}$ consists of only the ground-truth reward, the objective (8) is equivalent to standard RLHF (1), up to a constant offset independent of $\theta$. More generally, whenever $\mathcal{S}$ includes a good proxy $\widetilde{r}$ for $r^{*}$, the pessimistic advantage evaluation ensures that the policy $\pi^{*}_{\theta}$ that maximizes eq. 8 still has a large advantage over $\pi_{\mathrm{ref}}$ under all $r\in\mathcal{S}$, including $\widetilde{r}$. This use of pessimism to handle uncertainty in the knowledge of the true reward is related to similar techniques in the offline RL literature (Kumar et al., 2020; Cheng et al., 2022).
For the objective to be meaningful, however, the set $\mathcal{S}$ has to be chosen carefully. When $\mathcal{S}$ is small, it might not include any good proxy for $r^{*}$. Conversely, if $\mathcal{S}$ is too rich, it forces $\pi_{\theta^{*}}$ to be nearly identical to $\pi_{\mathrm{ref}}$, since any deviations from $\pi_{\mathrm{ref}}$ might be penalized by some reward model in $\mathcal{S}$. Consequently, we want to design $\mathcal{S}$ to be the smallest possible set which contains a reasonable approximation to $r^{*}$.
Finally, to solve (8), we can reformulate it as an equivalent constrained offline optimization problem, which conveniently admits a similar loss form as (7), as shown below:
@@ 5.2.1 Optimization
The constraint in (9) can be relaxed and approximately optimized by introducing an objective with a Lagrangian-style penalty with strength $\alpha>0$ on a form of distillation loss as (7), i.e.,
@@
Assume the same conditions as Theorem 1. Then for any $0<\gamma<\infty$, there exists a $\lambda\geq 0$ such that $\pi_{\theta^{*}}(y\mid x)\in\arg\!\min_{\pi_{\theta}}\mathcal{L}_{\mathrm{pdistill}}(\mathcal{S},\pi_{\theta};\rho)$, where $\pi_{\theta^{*}}$ is a minimizer over all possible policies of the objective (9), for the effective reward model set
@@
Takeaways: Pessimistic reward model distillation It can often be unclear what reward model should be used as a target for policy distillation. A natural optimization objective in the face of uncertainty is to optimize for the worst-case reward, out of a set of plausible reward candidates. We show that this can be formulated in a very similar style to the original distillation loss in Eq. (7) by simply adding an additional forward KL penalty term.
@@ 5.3 Pessimistic DPO
Proposition 4 can also be leveraged to obtain an alternative, implicitly pessimistic, objective that uses DPO directly instead of distillation. Consider the following regularized DPO loss:
@@ 6 Experimental results
The main motivation for reward distillation and pessimism is to increase alignment robustness in challenging settings where it is difficult to learn good policies directly from the preference data. To demonstrate the effectiveness of our approach, we run experiments on the popular TL;DR summarization task Stiennon et al. (2020); Völske et al. (2017), in which we simulate a scenario where the preference data has a spurious correlation between the length of a summary and whether or not it is preferred. [^6] Length has been repeatedly shown in the past to correlate with reward Singhal et al. (2023); Park et al. (2024). Additionally, we show results for an unbiased setting on TL;DR, as well for an unbiased setting on Anthropic Helpfulness Bai et al. (2022).
@@ 6.1 Experimental setup
We first train an “oracle” reward model on the TL;DR preference data training set Stiennon et al. (2020) and relabel all preference pairs with this oracle. This enables us to use the oracle reward model for evaluation, without worrying about the gap to true human preferences. After relabeling, longer responses (where longer is defined as $y_{1}$ having at least $10\%$ more tokens than $y_{2}$) are preferred in $61\%$ of the examples.
To test the effect of a spurious correlation on preference-based policy optimization, we select a training set of 30K examples from the relabeled data such that the longer output is preferred in $\rho$ fraction of examples, with $\rho\in\{0.2,0.3,0.4,0.5,0.6,0.7,0.8\}.$ Each such training set is denoted $\mathcal{D}_{\rho}$. At each $\mathcal{D}_{\rho}$, we compare our approach to DPO Rafailov et al. (2023) and IPO Azar et al. (2024), which are currently the most commonly used offline alignment methods. We test the following variants of distillation and pessimism:
@@
  • Pessimistic Distilled DPO (dp-DPO): Combines the above two by training a reward model $r_{\rho}$ on $\mathcal{D}_{\rho}$ and optimizing the pessimistic distillation objective (Eq. (11)) with confidence set $\mathcal{S}=\{r_{\mathrm{tgt}}\}$.
  • Pessimistic Ensemble DPO (e-DPO): To create ensembles of reward models, we subsample from each $\mathcal{D}_{\rho}$ five preference datasets, $\mathcal{D}_{\rho,b}$, at $b\in\mathcal{B}=\{0.2,0.4,0.5,0.6,0.8\}$, such that the fraction of pairs where the longer response is preferred is $b$, and train reward models $r_{\rho,b}$ on those subsets. Consequently, sensitivity to length should vary across ensemble members. We then apply the same procedure as dp-DPO above, with a confidence set $\mathcal{S}_{\rho}=\{r_{\rho,b}\}_{b=1}^{\mathcal{B}}$.
@@ 6.2 Results
Figure 2: Main results, showing the oracle reward compared to the initial finetuned policy (the oracle reward of the initial finetuned policy is $\approx-1$). Error bars correspond to bootstrap 95% confidence intervals for finite sample variance. Ensemble DPO (e-DPO) is significantly better than DPO and IPO in the challenging setup where shorter responses are preferred ($\rho\leq 0.5$), and is generally the best-performing method overall in this regime. Distilled DPO (d-DPO) performs best when longer responses are preferred ($\rho>0.6$).
We present the results of our experiment in Figure 2. As can be seen in the plot, the more challenging setting is when $\rho<0.5$, which corresponds to a sample of preference annotations in which shorter outputs are generally preferred. This distribution shift is more difficult because as mentioned the oracle reward model (trained on human annotations) has a bias in favor of longer outputs Singhal et al. (2023). Nevertheless we get sizable improvements compared to the reference policy $\pi_{\textrm{SFT}}$ for all length bias values.
@@
6.3 Additional results in an unbiased setting
To test the ability of our method to perform well on preference tasks where no bias is present, we next run experiments on the Anthropic Helpfulness dataset Bai et al. (2022). We use a Gemini 1.0 Ultra Gemini Team (2024) LLM-as-a-judge model for evaluating win-rates of the policies over both the SFT starting point and the best DPO baseline. As shown in Table 1, in this unbiased setting our distillation objectives can also provide modest gains. Concretely, e-DPO’s win rate against the SFT policy is 65.8%, while DPO’s win rate is 64.2%. Moreover, comparing e-DPO and DPO directly, e-DPO wins in 49.7% of the cases, while DPO wins in 46.9% of the cases (the rest are considered to be ties with no preference relation).
Takeaways: Experimental results Reward model distillation, specifically pessimistic reward model distillation when an ensemble of reward models is used, leads to improved robustness on tasks where bias is present in the preference data. In addition, reward model distillation results in small improvements in performance even on unbiased settings, making it a simple but compelling algorithmic modification to offline training.
Method %Wins %Ties %Losses
e-DPO vs. SFT 65.8 1.3 32.9
d-DPO vs. SFT 65.6 1.0 33.3
DPO vs. SFT 64.2 1.0 34.8
Method %Wins %Ties %Losses
e-DPO vs. DPO 49.7 3.4 46.9
d-DPO vs. DPO 48.2 3.6 48.2
Table 1: Side-by-side win rates on the Helpfulness dataset (with a Gemini 1.0 Ultra evaluator).
@@ 7 Theoretical analysis
This section characterizes solutions offered by pessimistic DPO and distillation to the issues identified in §4, focusing on the simplified scenario in which we optimize with respect to a single preference pair $(y^{w},y^{\ell})$.
7.1 Optima
In its Lagrangian formulation, pessimistic DPO adds a forward KL term to the DPO objective (§5.3). Here we seek to better analyze how this additional term effects the optimal policy. For the sake of analysis, we assume that the preference annotations are sampled from the reference distribution, $\mu(x)\times\pi_{\mathrm{ref}}(y\mid x)\times\pi_{\mathrm{ref}}(y\mid x)$. Then a finite-sample approximation of the forward KL term is
$\hat{\Omega}(\Theta):=\sum_{(y^{w},y^{\ell})\in\mathcal{D}_{\mathrm{Pref}}}-(\log\pi_{\theta}(y^{\ell})+\log\pi_{\theta}(y^{w})).$
By applying this finite-sample approximation, p-DPO has a finite optimum, unlike DPO, as shown in Proposition 1. Note that this analysis is limited in two ways: (1) as mentioned, we compute the KL term over the completions in the preference data; (2) we directly optimize the probability ratios $\psi_{w}=\pi_{\theta}(y^{w})/\pi_{\mathrm{ref}}(y^{w})$ and $\psi_{\ell}=\pi_{\theta}(y^{\ell})/\pi_{\mathrm{ref}}(y^{\ell})$, rather than optimizing them jointly through the parameters. For sufficiently expressive $\pi_{\theta}$, however, this approximation captures the behavior of the two algorithms reasonably well.